40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5937     Entry time: Fri Nov 18 00:36:23 2011
Author: Zach 
Type: Update 
Category: Green Locking 
Subject: Y-Arm PDH box modified 

I modified the Y-Arm PDH box (S/N 17) to have the same TF as the one of the temporary setup described in Kiwamu's earlier entryNote that the TF below was taken with the gain knob set to 0, so that the proper DC gain is achieved with a setting of ~4. This is desirable because it gives us wiggle room.

The changes were:

  • R14: 25 -> 50
  • R29: 1k -> 10.5k
  • R30: 1k -> 20k
  • R28: 102 -> OMIT
  • C20: 84nF -> OMIT
  • R31: SHORT -> 475
  • R16: 10k -> 48.7k
  • R24: 10 -> 5

Below is the TF along with the LISO model. They are different at low frequencies because the box must have been railing internally (though the phase shows that the result is as expected), and there is a feature around 60 kHz that probably arises from some op amp instability. I will see if adding a small cap somewhere does the trick, and then take a new TF with a lower source voltage.

pdh17_tf_vs_liso_11_17_11.png

I'll try to lock the arm with the box tomorrow.

ELOG V3.1.3-